# Maud Biquard ISAE-Supaero / CNES

### v1

@@ Abstract
Satellite optical images, upon their on-ground receipt, offer a distorted view of the observed scene. Their restoration, classically including denoising, deblurring, and sometimes super-resolution, is required before their exploitation. Moreover, quantifying the uncertainty related to this restoration could be valuable by lowering the risk of hallucination and avoiding propagating these biases in downstream applications. Deep learning methods are now state-of-the-art for satellite image restoration. However, they require to train a specific network for each sensor and they do not provide the associated uncertainties. This paper proposes a generic method involving a single network to restore images from several sensors and a scalable way to derive the uncertainties. We focus on deep regularization (DR) methods, which learn a deep prior on target images before plugging it into a model-based optimization scheme. First, we introduce VBLE-xz, which solves the inverse problem in the latent space of a variational compressive autoencoder, estimating the uncertainty jointly in the latent and in the image spaces. It enables scalable posterior sampling with relevant and calibrated uncertainties. Second, we propose the denoiser-based method SatDPIR, adapted from DPIR [73], which efficiently computes accurate point estimates. We conduct a comprehensive set of experiments on very high resolution simulated and real Pléiades images, asserting both the performance and robustness of the proposed methods. VBLE-xz and SatDPIR achieve state-of-the-art results compared to direct inversion methods. In particular, VBLE-xz is a scalable method to get realistic posterior samples and accurate uncertainties, while SatDPIR represents a compelling alternative to direct inversion methods when uncertainty quantification is not required.
@@ 1 Introduction
Satellite imaging is of considerable interest for various remote sensing applications such as environmental monitoring, conservation programs, or urban planning. Some satellite optical systems can now acquire images at a very high resolution. For instance, a PLEIADES-HR satellite captures images in which each pixel represents a $70\text{cm}\times 70$cm area on the ground and is coded in 12 bits. This allows for a detailed visualization of the Earth surface and provides extremely valuable information for object detection [27, 10], change detection, or semantic segmentation [69, 4] . However, despite constant technological advances in sensor accuracy, satellite images provide an imperfect representation of the observed scene. In particular, the images contain noise and blur induced by the optical system, satellite movement, atmospheric perturbations, and onboard compression. As a consequence, the received image must be restored to get the most out of its on-ground exploitation. The restoration aims at enhancing the image quality by removing the noise and blur mentioned above. Additional processing, such as super-resolution, can also be applied to ease data interpretation for downstream tasks. All these processes must preserve the integrity of the information: no information should be removed and no false details added. Furthermore, the restoration process has to remain computationally viable to treat the important amount of data transmitted on the ground. Finally, the restoration process should ideally be sufficiently generic to process images from different satellites.
@@
We restrict our study to panchromatic images, which offer the highest spatial resolution. They can be combined with multi-spectral images of lower resolution through pansharpening to obtain high resolution multispectral images. The acquisition of a panchromatic image by the satellite optical system can be modelled as
$y=\mathcal{D}_{0}(h_{0}*s)+w,$ (1)
@@
where $s$ represents the observed landscape, $y$ the acquired image, $h_{0}=h_{atmo}*h_{move}*h_{ins}$ models the combined effects of the atmosphere, of the movement during integration time, and of the instrument. $\mathcal{D}_{0}$ denotes the sampling operator. Vector $w$ represents a white Poisson-Gaussian noise which can, for the considered problems, be approximated by a Gaussian noise of variance $\sigma_{w}^{2}(s)=\sigma_{0}^{2}+K\mathcal{D}_{0}(h_{0}*s)$, where $\sigma_{0}$ and $K$ are noise parameters specific to a given optical system. Current Pléiades image restoration processes [40] employ traditional methods such as NL-Bayes for denoising [42] and inverse filtering for deconvolution [40]. However, these methods require ad hoc operations to stick to the white Gaussian noise model favorable to NL-Bayes denoiser [49]. Moreover, since the advent of deep learning, they are not anymore state-of-the-art for image restoration.
@@
Indeed, deep learning has led to a substantial breakthrough in image restoration [73, 79]. A first category of methods, which we denote by direct inversion methods in the sequel, learns a neural network that maps the degraded image to the restored image in a supervised manner [79, 68, 66]. However, as they learn a specific mapping for each forward model defined by (1), a neural network must be learned for each sensor. A second category of methods, which we call deep regularization (DR) methods in the following, merely learns a regularization independently of the forward model (1) from a dataset of target images, that is the original landscapes downsampled at a target resolution. While plug-and-play (PnP) approaches learn an implicit regularization within a denoiser [73, 64, 31, 34], latent optimization methods regularize explicitly the inverse problem, by looking for the problem solution in the latent space of a generative model [6, 5, 56]. DR methods typically compute the Maximum A Posteriori (MAP) estimate:
@@
where $x$ is the current estimation of the target image, $y$ is the measured image, $p_{Y|X}(y|x)$ is the data likelihood, $p_{X|Y}(x|y)$ the inverse problem posterior. The regularization $-\log p_{X}(x)$ promotes solutions of the inverse problem most compatible with the prior $p_{X}(x)$, which is learnt beforehand from a dataset of target images. A first benefit is that a single neural network is trained once to perform different restoration tasks for different sensors or different acquisition conditions, in the spirit of foundation models. A second benefit is that DR methods, unlike direct inversion approaches, can use standard denoisers or generative models, for which models pretrained on large databases are available. Hence, using them for satellite restoration only requires their finetuning on a smaller database of target satellite images. Then, DR methods rely on an optimization process often similar to eq. 2, which allows to control the strength of the regularization during the optimization process, for instance with a classical regularization parameter. This helps to remain faithful to the measure, which is crucial for satellite image restoration.
@@
Last but not least, some DR methods do not only estimate the MAP, but enable to sample from the posterior distribution $p_{X|Y}(x|y)$. This remarkably allows to derive uncertainties, which are valuable for downstream applications. However, these methods are often computationally expensive as they mainly rely on Monte Carlo Markov Chains (MCMC) sampling. Yet, in Biquard et al. [5], we have proposed another method for approximate posterior sampling, called Variational Bayes Latent Estimation (VBLE), relying on latent optimization with a variational compressive autoencoder. VBLE allows for approximate posterior sampling with no significant increase in computational cost, and seems well suited for remote sensing image restoration.
However, while direct inversion methods are more and more studied in the context of remote sensing, and even used in real on-ground segments  [81, 12], DR methods have rarely been considered in the literature. This paper aims at filling this gap, by demonstrating the interest of DR for remote sensing image restoration. In particular, we will show that DR methods can reach similar or even better performance than inversion-based techniques, and with the extra ability to predict an accurate estimation of the uncertainties. Although they require in general more computation, this remains scalable in a real context. And since they can be easily adapted to any sensor or acquisition conditions without any re-training, we believe they can even be easier to use in real on-ground segments.
Our contributions are the following:
  • We propose an improved version of VBLE, called VBLE-xz, which yields more informative uncertainties by taking into account the reconstruction error of the generative model. We also adopt a calibration strategy, so that our method provides state-of-the-art results with accurate perpixel error bounds in a reasonable amount of time.
  • We also propose SatDPIR, an adaptation of the state-of-the-art PnP method Deep Plug-and-play Image Restoration (DPIR) [73] to the specific noise model encountered in satellite imaging. SatDPIR allows for fast and accurate satellite image restoration and can be advantageously used when uncertainties are not required.
@@
  • We conduct a comprehensive set of experiments on high resolution satellite images simulated in a realistic manner, as well as on real satellite images to show the interest of these two DR methods. We show their interest over alternative DR techniques. Furthermore, we also compare our methods to direct inversion in terms of performance and robustness, showing that they yield similar metrics as direct inversion methods, but with fewer artefacts and improved robustness to modelling errors.
@@ 2.1 Image restoration techniques
Image restoration is a well-known inverse problem, which has classically been addressed using variational methods. The key point of these methods, which consist in minimizing a functional composed of a data fidelity term and of a regularization term such as in eq. 2, is to choose a proper regularization. Classical regularizations include total variation [57], Tikhonov regularization [63], and sparsity-promoting penalties on well-chosen representations such as wavelet bases or dictionaries [62, 21, 50]. Recently, deep learning based methods have emerged as a powerful alternative to these classical methods. Direct inversion methods, which directly learn a mapping from the degraded image to the restored image using a neural network, are widespread in image restoration. They commonly use convolution neural networks (CNNs) within either a U-Net architecture [68, 71, 48], or a ResNet architecture [74, 43, 66]. Generative Adversarial Networks (GANs) have been also used to improve perceptual performance [43, 66]. Recently, transformer-based method have enabled better results [68, 70, 44], as well as diffusion-based methods [47, 48], yet at the cost of an increased computational complexity.
@@
Unlike direct inversion methods, DR methods learn only the regularization term, with the help of denoisers or diffusion models in PnP methods [64], or with generative models such as Variational Autoencoders (VAEs) [38] or GANs [26] for latent optimization methods. PnP methods, which yield state-of-the-art results on a wide variety of inverse problems, consist in using splitting algorithms, such as Alternating Direction Method of Multipliers (ADMM) [64, 9] or Half Quadratic Splitting (HQS) [73, 31], that separately handle the data term and the regularization term in the optimization problem. Their key concept is to use Gaussian denoisers [35], and in particular deep denoisers [75] in place of the proximal operator of the regularization. Recently, the use of denoising diffusion models [28] has led to remarkable results for solving imaging inverse problems [36, 13, 82].
@@
While PnP methods implicitly learn the regularization through the denoising process, latent optimization methods represent a more direct way to regularize inverse problems by estimating the data distribution within generative models. The seminal work [6] seeks the inverse problem solution in the latent space of a VAE or a GAN, by optimizing a functional, leading to a latent MAP estimate. Then, the solution is taken as the image generated from the latent MAP. This approach is interesting because of its explicit formulation as well as its well-defined Bayesian framework. However, it suffers from its dependency to the quality of the generative model since the obtained solution necessarily lies in the generator range. To tackle this problem, [15, 14] permit small deviations from the generative manifold while [25, 18, 19] propose to jointly optimize $x$ and $z$ using splitting algorithms. More recently, in [5], we proposed to replace the generative model by a compressive autoencoder, which can generate a wide variety of images while efficiently regularizing the inverse problem in the latent space through its hyperprior.
@@
While most of the methods described above only provide a single point estimate to the inverse problem, many applications, particularly in Earth observation, require an estimation of the confidence in this solution. To this end, several Bayesian methods enable to sample the posterior distribution $p_{X|Y}(x|y)$ of the inverse problem solution. Some methods yield stochastic solutions to the inverse problem, relying, for instance, on the implicit prior provided by Gaussian denoisers [34] or exploiting diffusion model properties [82]. Another type of method uses Markov Chain Monte Carlo (MCMC) to sample from the true posterior. In particular, [20] proposes to use Unadjusted Langevin Algorithm (ULA) to solve imaging inverse problems. Building on this, PnP-ULA [41] approximates the log-likelihood gradient in ULA by employing a Gaussian denoiser in a PnP framework. Interestingly, [29] attempts to combine MCMC with latent optimization methods by designing a MCMC sampling scheme in the latent space of a generative model. These approaches enable to sample from the true posterior but require a lot of iterations to converge, making them hardly scalable for satellite imaging. In [5], we introduced the VBLE algorithm, performing variational inference in the latent space of a compressive autoencoder. With negligible additional computation cost compared to the deterministic algorithm, VBLE represents a relevant alternative to MCMCs for posterior sampling.
@@ 2.2 Application to satellite optical imaging
Once transmitted to the ground station and decompressed, satellite images are subject to an in-depth semantic analysis including for instance perpixel classification [69, 4], object detection [27, 10], or 3D reconstruction [22]. Previously, satellite images must be restored and eventually super-resolved. Most of the deep learning currently used methods are based on direct inversion. For satellite single image super-resolution (SISR) tasks, CNN-based methods [72, 45, 55, 23] and GAN-based methods [7, 33] are widely used. In particular, [81, 12] employ a SRResNet architecture for joint denoising, deblurring, and super-resolution of optical satellite images. Furthermore, [30] proposes a deep pansharpening network, and [77, 80] perform deep hyperspectral super-resolution.
Concerning DR methods, PnP methods are commonly used for hyperspectral image processing tasks. For instance, [65, 39] use ADMM PnP framework for hyperspectral image restoration, and [61] for hyperspectral image sharpening. To process panchromatic or multispectral satellite images, [46] employs ADMM for pansharpening, while [60] employs DPSR algorithm [76] for super-resolution. Up to our knowledge, PnP methods have not been commonly used for high resolution satellite image restoration. Concerning latent optimization methods, they have been successfully applied to medical imaging [19]. However, due to the complexity of using these methods on complex datasets, they have not been yet applied to satellite imaging.
@@
3 SatDPIR: An adapted plug-and-play algorithm for satellite image restoration
3.1 Background on DPIR
DPIR [73] is a PnP algorithm relying on Half Quadratic Splitting (HQS). Its particularly low number of iterations as well as its convenient parameter tuning have made DPIR a reference PnP method for solving inverse problems with white Gaussian noise.
Considering a forward model $y=Ax+w$ with $A$ the degradation operator, for instance a convolution, and a white Gaussian noise $w\sim\mathcal{N}(0,\sigma^{2})$, DPIR seeks to minimize $\frac{1}{2\sigma^{2}}||y-Ax||^{2}+\lambda R(x)$ with respect to $x$ where $R$ is a regularization term. HQS algorithm consists of the following splitting:
$\min_{x,u}\frac{1}{2\sigma^{2}}||y-Ax||^{2}+\lambda R(u)+\frac{\mu}{2}||u-x||^{2},$ (3)
with $u$ an auxiliary variable that is jointly optimized with $x$. This leads to an alternate optimization scheme
$x_{k}$ $=\operatorname*{arg\,min}_{x}\frac{1}{2\sigma^{2}}||y-Ax||^{2}+\frac{\mu}{2}||x-u_{k-1}||^{2},$ (4)
$u_{k}$ $=\operatorname*{arg\,min}_{u}\frac{1}{2\sqrt{\lambda/\mu}^{2}}||u-x_{k}||^{2}+R(u).$ (5)
Equation 4 is a proximal step on the data fidelity which can be computed in closed form for various inverse problems. Equation 5 amounts to a denoising step with noise level $\sigma_{d}=\sqrt{\lambda/\mu}$. [73] proposes to use decreasing values of $\sigma_{d}$, evenly spaced in log scale from $\sigma_{1}$ to $\sigma_{2}<\sigma_{1}$ over the iterations. $\sigma_{2}=\sigma$ corresponds to the noise level in the measurement, while $\sigma_{1}>\sigma$ is a hyperparameter that is typically set to $50/255$ for 8 bits images. With $\lambda$ a fixed regularization parameter, this leads to a larger $\mu=\frac{\lambda}{\sigma_{d}^{2}}$ at each iteration. Note that the number of iterations is fixed a priori because of the noise schedule. This results in a very fast convergence, generally within 8 iterations. In [73], the successive denoising steps are performed with DRUnet denoiser as it can manage different noise levels.
3.2 Proposed DPIR adaptation
DPIR algorithm has been specifically designed for inverse problems with white Gaussian noise of fixed variance. Yet, in the forward model (1) considered for satellite image restoration, the acquisition noise is Gaussian but with a spatially varying variance $\sigma_{w}^{2}(x)=\sigma_{0}^{2}+K(h_{0}*x)$. The associated data likelihood substantially complicates the problem resolution. In particular, the proximal step of eq. 4 can no longer be computed in closed form. This significantly slows down the algorithm convergence for large images, such as satellite ones. To overcome this limitation, we propose in this section a smart initialization of the proximal gradient descent that considerably speeds up its convergence.
First, for the considered inverse problem, the quadratic data fidelity term in equations (3) and (4) should be replaced by the real negative log-likelihood:
$-\log p_{Y|X}(y|x)\propto\frac{1}{2}(y-h_{0}*x)^{T}$ $\Sigma_{w}(x)^{-1}(y-h_{0}*x)$
$+\frac{1}{2}\log|\Sigma_{w}(x)|$ (6)
up to a constant, with $\Sigma_{w}(x)=\operatorname{diag}(\sigma_{w}^{2}(x))$. Thus, we replace eq. 4 with:
$x_{k}$ $=\arg\min_{x}-\log p_{Y|X}(y|x)+\frac{\mu}{2}||x-u_{k-1}||^{2}.$ (7)
We adapt the noise schedule by considering $\sigma_{2}=\sqrt{a}$, that is the minimum level of noise in the image, and $\sigma_{1}$ is empirically reduced to $20/255$, which corresponds to $320/4095$ for 12-bit images, as the considered noise levels for satellite imaging is much smaller than classically considered noise levels.
Then, eq. 7 can no longer be expressed in closed form, and should be computed using gradient descent. We propose to initialize the gradient descent at iteration $k$ with
$x_{k}^{(0)}=\arg\min_{x}\frac{1}{2\bar{\sigma}^{2}}||y-h_{0}*x||^{2}+\frac{\mu}{2}||x-u_{k-1}||^{2}$ (8)
where $\bar{\sigma}^{2}=a+b(h_{0}*\bar{u}_{k-1})$ and $\bar{u}_{k-1}$ is the mean value of $u_{k-1}$. Equation 8, which corresponds to the proximal operator considering a fixed noise level $\bar{\sigma}^{2}$, can be computed in closed form. Then, we consider two phases during DPIR iterations. During the first half of the iterations, where each iteration corresponds to significant changes in $x_{k}$, we do not perform gradient descent on eq. 8 but only the proximal step eq. 8, that is $x_{k-1}=x_{k-1}^{(0)}$. Indeed, we consider $x_{k}^{(0)}$ to be a sufficiently good approximation to $x_{k}$ in this case. During the second half of the iterations, we perform gradient descent on eq. 7 starting from $x_{k}^{(0)}$ during only 5 iterations. We show in the experiments section that this optimization strategy produces very similar results as the original process while being much faster.

### v2

@@ Abstract
Satellite optical images, upon their on-ground receipt, offer a distorted view of the observed scene. Their restoration, including denoising, deblurring, and sometimes super-resolution, is required before their exploitation. Moreover, quantifying the uncertainties related to this restoration helps to reduce the risks of misinterpreting the image content. Deep learning methods are now state-of-the-art for satellite image restoration. Among them, direct inversion methods train a specific network for each sensor, and generally provide a point estimation of the restored image without the associated uncertainties. Alternatively, deep regularization (DR) methods learn a deep prior on target images before plugging it, as the regularization term, into a model-based optimization scheme. This allows for restoring images from several sensors with a single network and possibly for estimating associated uncertainties. In this paper, we introduce VBLE-xz, a DR method that solves the inverse problem in the latent space of a variational compressive autoencoder (CAE). We adapt the regularization strength by modulating the bitrate of the trained CAE with a training-free approach. Then, VBLE-xz estimates relevant uncertainties jointly in the latent and in the image spaces by sampling an explicit posterior estimated within variational inference. This enables fast posterior sampling, unlike state-of-the-art DR methods that use Markov chains or diffusion-based approaches. We conduct a comprehensive set of experiments on very high-resolution simulated and real Pléiades images, asserting the performance, robustness and scalability of the proposed method. They demonstrate that VBLE-xz represents a compelling alternative to direct inversion methods when uncertainty quantification is required. The code associated to this paper is available in https://github.com/MaudBqrd/VBLExz.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
Index terms— Satellite image restoration, Deep regularization, Latent optimization, Uncertainty quantification, Posterior sampling, Plug-and-play methods
@@ 1 Introduction
Satellite imaging is of considerable interest for various remote sensing applications such as environmental monitoring, conservation programs, or urban planning. Some satellite optical systems can now acquire images at a very high resolution. For instance, a PLEIADES-HR satellite captures images in which each pixel represents a $70\text{cm}\times 70$cm area on the ground and is coded in 12 bits. This allows for a detailed visualization of the Earth’s surface and provides extremely valuable information for object detection [Han2014, Chen2014], change detection, or semantic segmentation [Yang2018, Benediktsson2005]. However, despite constant technological advances in sensor accuracy, satellite images provide an imperfect representation of the observed scene. In particular, the images contain noise and blur induced by the optical system, satellite movement, atmospheric perturbations, and onboard compression. As a consequence, the received image must be restored to get the most out of its on-ground exploitation. The restoration aims at enhancing the image quality by removing the noise and blur mentioned above and may also be associated with super-resolution or inpainting that ease downstream data interpretation. All these processes must preserve the integrity of the information: no information should be removed and no false details added. Furthermore, the restoration process has to remain computationally viable to treat the important amount of data transmitted on the ground. Finally, the restoration process should ideally be sufficiently generic to process images from different sensors.
@@
We restrict our study to panchromatic images, which offer the highest spatial resolution. They can be combined with multi-spectral images of lower resolution through pansharpening to obtain high-resolution multispectral images. The acquisition of a panchromatic image by the satellite optical system can be modeled as
$y=\mathcal{D}_{0}(h_{0}*l)+w,$ (1)
@@
where $l$ represents the observed landscape, $y$ the acquired image, $h_{0}=h_{atmo}*h_{move}*h_{ins}$ models the combined effects of the atmosphere, of the movement during integration time, and of the instrument. $\mathcal{D}_{0}$ denotes the sampling operator. Vector $w$ represents a white Poisson-Gaussian noise which can, for the considered problems, be approximated by a Gaussian noise of variance $\sigma_{w}^{2}(l)=\sigma_{0}^{2}+K\mathcal{D}_{0}(h_{0}*l)$, where $\sigma_{0}$ and $K$ are noise parameters that are specific to a given optical system. Current Pléiades image restoration processes [Latry2012] employ traditional methods such as NL-Bayes for denoising [Lebrun2013] and inverse filtering for deconvolution [Latry2012]. However, these methods require ad hoc operations to stick to the white Gaussian noise model favorable to NL-Bayes denoiser [Makitalo2012]. Moreover, since the advent of deep learning, they are not state-of-the-art for image restoration anymore.
@@
Indeed, deep learning has led to a substantial breakthrough in image restoration [Zhang2021, Zhang2018a]. A first category of methods, which we denote by direct inversion methods in the sequel, learns a neural network that maps the degraded image to the restored image in a supervised manner [Zhang2018a, Wang2022, Wang2018]. However, as they learn a specific mapping for each forward model defined by (1), a neural network must be learned for each sensor. A second category of methods, which we call deep regularization (DR) methods in the following, merely learns a regularization independently of the forward model (1) from a dataset of target images, that is the original landscapes sampled at a target resolution. While plug-and-play (PnP) approaches learn an implicit regularization within a denoiser [Zhang2021, Venkatakrishnan2013, Hurault2022, Kadkhodaie2021], latent optimization methods regularize explicitly the inverse problem, by looking for the problem solution in the latent space of a generative model [Bora2017, Biquard2023, Prost2023]. DR methods typically compute the Maximum A Posteriori (MAP) estimate:
@@
where $x$ is the current estimation of the target image, $y$ is the measured image, $p_{Y|X}(y|x)$ is the data likelihood, $p_{X|Y}(x|y)$ the inverse problem posterior. The regularization $\log p_{X}(x)$ promotes solutions of the inverse problem most compatible with the prior $p_{X}(x)$, which is learned beforehand from the dataset of target images.
@@
Last but not least, some DR methods do not only estimate the MAP, but enable to sample from the posterior distribution $p_{X|Y}(x|y)$. This remarkably enables to derive uncertainties, which are valuable for downstream applications. However, these methods are often computationally expensive as they mainly rely on Markov Chain Monte Carlo (MCMC). To overcome this, in Biquard et al. [Biquard2023], we have proposed another method for approximate posterior sampling, called Variational Bayes Latent Estimation (VBLE). By approximating the inverse problem posterior in the latent space of a variational compressive autoencoder (CAE), VBLE allows for approximate posterior sampling with no significant increase in computational cost and seems well-suited for remote sensing image restoration. However, VBLE possesses two remaining drawbacks. First, it requires training CAEs at different bitrates to adapt the regularization strength to the inverse problem difficulty. Second, the modeling of the posterior distribution in the latent space fails to fully account for all uncertainties, such as the representation error between the true solution and its projection onto the range of the generative model.
In this paper, we introduce VBLE-xz, a new method derived from VBLE that addresses those two limitations. Besides, while direct inversion methods are more and more studied in the context of remote sensing, and currently used in on-ground segments [Zhu2020, Chouteau2022], DR methods have been rarely considered in the associated literature. We show here that VBLE-xz, and more generally DR methods, yield numerous advantages compared to them. Precisely, our contributions are the following:
  • We propose an improved version of VBLE, called VBLE-xz. In particular, VBLE-xz models the posterior distribution of the inverse problem jointly in the latent and in image space, leading to an improved prediction of the uncertainties by a large margin.
  • To restore inverse problems of diverse difficulties with the same network, we combine VBLE-xz with a multirate CAE. Following the procedure proposed in [Arezki2024], a single CAE is trained once, at a high bitrate. Then, this bitrate can be simply modified, during inference and without any finetuning. We adapt this framework to satellite image restoration and we show that it adjusts remarkably well to the inverse problem difficulty.
@@
  • We conduct a comprehensive set of experiments on high-resolution satellite images simulated in a realistic manner and on real satellite images. These experiments demonstrate two important points. First, we show that VBLE-xz yields state-of-the-art results in terms of restoration performance, uncertainty quantification, and computational load compared to other recent posterior sampling methods. Second, we also compare the DR methods VBLE-xz and PG-DPIR [Biquard2025] to direct inversion in terms of performance and robustness, showing that they remain competitive in terms of metrics, but with fewer artifacts and improved robustness to modeling errors. This demonstrates that deep regularization is a compelling alternative to direct inversion for real satellite processing chains.
@@ 2.1 Image restoration techniques
Image restoration is a well-known inverse problem, which has classically been addressed using variational methods. The key point of these methods, which consist in minimizing a functional composed of a data fidelity term and a regularization term such as in eq. 2, is to choose a proper regularization. Classical regularizations include total variation [Rudin1992], Tikhonov regularization [Tikhonov1963], and sparsity-promoting penalties on well-chosen representations such as wavelet bases or dictionaries [Tibshirani1996, Elad2010, Mallat2008]. Recently, deep learning-based methods have emerged as a powerful alternative to these classical methods. Direct inversion methods, which directly learn a mapping from the degraded image to the restored image using a neural network, are widespread in image restoration. They commonly use convolution neural networks (CNNs) within either a U-Net architecture [Wang2022, Zamir2021, Luo2023], or a ResNet architecture [Zhang2017, Ledig2017, Wang2018]. Generative Adversarial Networks (GANs) have also been used to improve perceptual performance [Ledig2017, Wang2018]. Recently, transformer-based methods have enabled better results [Wang2022, Zamir2022, Liang2021], as well as diffusion-based methods [Lugmayr2022, Luo2023], yet at the cost of an increased computational complexity. Most of the above mentioned methods assume known degraded-target image pairs, where the degraded images, often simulated, may not be totally realistic. Therefore, a domain gap may exist when applying those methods on real-world problems. Blind methods, such as [Lin2024, Xiao2025], aim to jointly restore the image and estimate the degradation model.
@@
Unlike direct inversion methods, DR methods learn only the regularization term, with the help of denoisers or diffusion models in PnP methods [Venkatakrishnan2013], or with generative models such as Variational Autoencoders (VAEs) [DiederikPKingma2014] or GANs [Goodfellow2014] for latent optimization methods. PnP methods, which yield state-of-the-art results on a wide variety of inverse problems, consist in using splitting algorithms, such as Alternating Direction Method of Multipliers (ADMM) [Venkatakrishnan2013, Chan2016] or Half Quadratic Splitting (HQS) [Zhang2021, Hurault2022], that separately handle the data term and the regularization term in the optimization problem. Their key concept is to use Gaussian denoisers [Kamilov2017], in particular deep denoisers [Zhang2017a] in place of the proximal operator of the regularization. Recently, the use of denoising diffusion models [Ho2020] has led to remarkable results for solving imaging inverse problems [Kawar2022, Chung2022, Zhu2023].
@@
While PnP methods implicitly learn the regularization through the denoising process, latent optimization methods represent a more direct way to regularize inverse problems by estimating the data distribution within generative models. The seminal work [Bora2017] seeks the inverse problem solution in the latent space of a VAE or of a GAN, by optimizing a functional, leading to a latent MAP estimate. Then, the solution is taken as the image generated from the latent MAP estimate. This approach is interesting because of its explicit formulation as well as its well-defined Bayesian framework. However, it suffers from its dependency on the quality of the generative model since the obtained solution necessarily lies in the generator range. To tackle this problem, [Dhar2018, Dean2020] permit small deviations from the generative manifold while [Gonzalez2022, Duff2022, Duff2023] propose to jointly optimize $x$ and $z$ using splitting algorithms. More recently, in [Biquard2023], we proposed to replace the generative model by a compressive autoencoder, which can generate a wide variety of images while efficiently regularizing the inverse problem in the latent space through its hyperprior.
@@
While most of the methods described above only provide a single point estimate of the inverse problem, many applications, particularly in Earth observation, require an estimation of the confidence in this solution. To this end, several Bayesian methods enable to sample the posterior distribution $p_{X|Y}(x|y)$ of the inverse problem solution. Some methods yield stochastic solutions to the inverse problem, relying, for instance, on the implicit prior provided by Gaussian denoisers [Kadkhodaie2021] or exploiting diffusion model properties [Zhu2023]. Another type of method uses Markov Chain Monte Carlo (MCMC) to sample from the true posterior. In particular, [Durmus2018] proposes to use Unadjusted Langevin Algorithm (ULA) to solve imaging inverse problems. Building on this, PnP-ULA [Laumont2022a] approximates the log-likelihood gradient in ULA by employing a Gaussian denoiser in a PnP framework. Interestingly, [Holden2022] attempts to combine MCMC with latent optimization methods by designing an MCMC sampling scheme in the latent space of a generative model. These approaches enable to sample from the true posterior but require a lot of iterations to converge, making them hardly scalable for satellite imaging. Alternatively, [Marinescu2020, Dasgupta2024, Kothari2021, Biquard2023] propose to learn an approximate posterior distribution in the latent space of a generative model using variational inference. These approaches are usually much faster than other posterior sampling methods. In particular, in [Biquard2023], we introduced the VBLE algorithm, performing variational inference in the latent space of a compressive autoencoder. With negligible additional computation cost compared to the deterministic algorithm, VBLE represents a relevant alternative to MCMCs for posterior sampling.
@@ 2.2 Application to satellite optical imaging
Once transmitted to the ground station and decompressed, satellite images are subject to an in-depth semantic analysis including for instance perpixel classification [Yang2018, Benediktsson2005], object detection [Han2014, Chen2014], or 3D reconstruction [Facciolo2017]. Previously, satellite images must be restored and eventually super-resolved. Most of the deep learning methods currently used are based on direct inversion. For satellite single image super-resolution (SISR) tasks, CNN-based methods [Zhang2013, Liebel2016, Pouliot2018, Fu2018] and GAN-based methods [Bosch2018, Jiang2019] are widely used. In particular, [Zhu2020, Chouteau2022] employ an SRResNet architecture for joint denoising, deblurring, and super-resolution of optical satellite images. Furthermore, [Huang2015] proposes a deep pansharpening network, and [Zhang2020, Zheng2021] perform deep hyperspectral super-resolution.
Concerning DR methods, PnP methods are commonly used for hyperspectral image processing tasks. For instance, [Wang2023, Lai2022] use ADMM PnP framework for hyperspectral image restoration, and [Teodoro2017] for hyperspectral image sharpening. To process panchromatic or multispectral satellite images, [Liu2022a] employs ADMM for pansharpening, while [Tao2020] employs DPSR algorithm [Zhang2019] for super-resolution. Up to our knowledge, PnP methods have not been commonly used for high-resolution satellite image restoration. Concerning latent optimization methods, they have been successfully applied to medical imaging [Duff2023]. However, due to the complexity of using these methods on less structured datasets, they have yet not been applied to satellite imaging.
@@
2.3 Position to related works
To our knowledge, this work is the first to evaluate DR methods on high-resolution satellite image restoration tasks. Most existing works in remote sensing rely on direct inversion approaches, as in [Chouteau2022, Zhu2020]. However, DR methods offer several advantages over direct inversion for remote sensing image restoration. In particular, direct inversion typically requires training a dedicated network for each sensor, along with a computationally intensive data simulation process, as realistic pairs of degraded and reference images must be generated under a variety of acquisition conditions. In contrast, DR methods enable a single neural network to be trained once and applied across multiple restoration tasks and sensors, aligning with the philosophy of foundation models [Bommasani2021]. Since this model is trained solely on target images, it significantly simplifies the data simulation pipeline. Furthermore, our work involves a careful design of realistic degraded and reference satellite images, generated from very high-resolution airborne images. This stands in contrast to many existing studies in remote sensing, where unrealistic degradation models are often used due to the scarcity of reference data [Liebel2016, Bosch2018, Jiang2019].
Then, VBLE-xz is a variational inference method. The key advantage of these methods for image restoration is that they enable very efficient posterior sampling by explicitly approximating the posterior. On the contrary, diffusion-based methods [Kawar2022, Chung2022, Zhu2023] require to run a diffusion process for each sample to draw, while MCMC methods [Holden2022, Laumont2022a] run a Markov chain requiring thousands of iterations to converge. This makes them computationally heavy and, in practice, hardly scalable for satellite image restoration pipelines.
Last, while variational inference has been commonly used for solving specific inverse problems [Tonolini2020, Gao2022] in a supervised or semi-supervised setting, some approaches [Kothari2021, Dasgupta2024, Marinescu2020, Biquard2023] propose to use it in a latent optimization setting. Precisely, in [Kothari2021, Dasgupta2024], small normalizing flows [Dinh2015] are used to approximate the posterior in the latent space of a normalizing flow or a GAN. Alternatively, [Marinescu2020, Biquard2023] are more closely related to VBLE-xz as they learn a simple unimodal approximate posterior in the latent space of a VAE or of a CAE. However, VBLE-xz significantly differs from these approaches. First, all the considered methods approximate the posterior distribution only in the latent space, while VBLE-xz approximates the joint latent and image posterior. Although [Gonzalez2022, Prost2023] also leverages this joint posterior to compute a deterministic Maximum A Posteriori, VBLE-xz approximates this joint posterior approximation using variational inference, which is completely novel. Second, the use of a multi-rate CAE [Arezki2024] to efficiently adapt the regularization strength enables to solve inverse problems of different difficulty with the same network. To our knowledge, such an approach to adapt the regularization has never been done for image restoration.